OpenStudy (anonymous):

slope -2/3, y intercept (0,5)

jimthompson5910 (jim_thompson5910):

start with y = mx + b, which is your general slope intercept form

jimthompson5910 (jim_thompson5910):

where m = slope b = y-intercept

jimthompson5910 (jim_thompson5910):

because the slope is -2/3, we know that m = -2/3 and because the y intercept is (0,5), we know that b = 5

jimthompson5910 (jim_thompson5910):

put this all together to go from y = mx+b to \[\large y = -\frac{2}{3}x + 5\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@jim_thompson5910 is right @frogula (:

OpenStudy (anonymous):

thank you i have problems doing these

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

ok well i may have another for you tho

OpenStudy (anonymous):

write each equation in slope-intercept form then identify the slope and the y -intercept for each line 2x-y=12

jimthompson5910 (jim_thompson5910):

solve 2x-y=12 for y in other words, isolate y

OpenStudy (anonymous):

so would it be 6 then

jimthompson5910 (jim_thompson5910):

2x-y=12 -y=12-2x ... Subtract 2x from both sides. y = -12 + 2x ... multiply everything by -1 to isolate y y = 2x - 12 .. rearrange terms

jimthompson5910 (jim_thompson5910):

now compare y = 2x - 12 with y = mx+b